A body moves with initial velocity of 30ms and accelerates uniformly until it attains a velocity of 80ms. It then continues at that velocity for same time and accelerates to rest

The total time taken for the journey is 40s and the total distance traveled is 2550m. If the time spent accelerating is 1/3 of travelling at constant velocity. Calculate the acceleration


1
Expert's answer
2021-06-01T10:50:21-0400
"s=0.5at^2+v(T-2t)+0.5at^2\\s=at^2+at(T-2t)\\2550=5^2a+5(30)a\\a=14.6\frac{m}{s^2}"